# Parallel Scan on Ascend AI Accelerators

### v1

@@ 3.1. Ascend Hardware
[Architecture of Ascend 910B]Architecture of Ascend 910B Training series accelerators. It depicts the so-called “split” architecture, i.e., cube and vector cores are split.
3.2. AscendC Programming Model
@@
4. Scan Algorithms on Ascend
Matrix multiplication is an essential operator in our discussions here; hence, we introduce some linear algebraic notation that will be used throughout the paper. We denote matrices using boldface font and capitals, i.e., $\bm{A},\bm{B},\bm{C}$. We use $s$ to denote the size of square matrices, i.e., $\bm{A}_{s}$, and define $\ell:=s^{2}$. We drop the subscript on the matrices when the dimension is clear from the context. We denote matrix multiplication between $\bm{A}$ and $\bm{B}$ by $\bm{C}:=\bm{A}\ @\ \bm{B}$. We denote by $\bm{U}_{s}$ the upper-triangular all-ones square matrix of size $s$, including ones on the main diagonal. $\bm{L}_{s}$ is the lower triangular all-ones of size $s$. $\bm{L}^{-}_{s}$ is the strictly lower triangular all-ones of size $s$: $\bm{L}_{s}^{-}$ has zeroes on the main diagonal. $\bm{1}_{s}$ denotes the all-ones square matrix of size $s$. We denote one-dimensional arrays using boldface font and lower-case: $\bm{r},\bm{x},\bm{y},\bm{z}$. We frequently partition an array $\bm{x}$ into tiles of length $\ell$, i.e., tiles are contiguous blocks of $\ell$ entries of $\bm{x}$. We note an arbitrary $\ell$-tile of $\bm{x}$ by $\bm{x}_{\ell}$. We also view a tile $\bm{x}_{\ell}$ as a row-major matrix $\bm{A}$ having $s$ rows and $s$ columns (pad with zeroes if necessary).
Next, in Section 4.2, we build on top of the ScanU and ScanUL1 algorithms and extend them into batched scan variants that operate on multi-dimensional arrays (tensors). Here, we discuss issues that arise when scheduling several scan operations into multiple cores, including padding, better load balancing on vector/cube units with a ratio of 2:1, etc.
Last but not least, in Section 4.3, we present a multi-core scan algorithm (MCScan, Algorithm 3) tailored for the Ascend AI accelerator. A key feature of MCScan is that it utilizes all the available cube and vector cores. MCScan is designed for scenarios involving very large one-dimensional arrays.
4.1. Warm-up: single cube scans
Figure 2. Diagram that shows the data path from an input tile $\bm{x}_{\ell}$ to an output tile $\bm{y}_{\ell}$ of the ScanU (Algorithm 1). Blue and green denote the input and output array, respectively.
[…]
[Algorithm 1)]Algorithm 1). Blue and green denote the input and output array, respectively.
The critical path or span of both proposed algorithms is linear on the input length for constant values of $s$, since there is a sequential dependency on the partial sums. Therefore, these kernels are more effective when the input array to be scanned has a relatively short length. Additionally, designing and developing scan algorithms that use a single cube core is a building block for extending these ideas to a multi-core scenario, as seen in Sections 4.2 and 4.3.
Figure 2 depicts the data movements and the memory view of ScanU. The input vector $\bm{x}$ is shown in blue. A tile $\bm{x}_{\ell}$ of $\bm{x}$ and $\bm{U}_{s}$ are loaded into the cube unit where the matrix multiplication occurs. The matrix multiplication result is written to global memory. The vector unit reads the cube unit output tile and propagates the prefix sum in place. The whole process, which consists of memory transfers and cube/vector operations, is pipelined over the input tiles using AscendC software pipeline capability.
1:procedure ScanU($\bm{x}$, $s$)
2:     Let $\bm{y}$ be the output array
3:     $\textit{partial}\leftarrow 0$triangleright Accumulation value (Vector Unit)
4:     Load $\bm{U}_{s}$ in $\mathtt{L0B}$
5:     for each $s^{2}$-tile of $\bm{x}$: $\bm{x}_{\ell}$ do $\triangleright$ Pipelined execution
6:         Load $\bm{x}_{\ell}$ to $\mathtt{L0A}$
7:         $\bm{C}\leftarrow\bm{A}_{s}\ @\ \bm{U}_{s}$triangleright acc. off, free $\mathtt{L0A}$
8:         Copy $\bm{C}$ from $\mathtt{L0C}$ to $\bm{y}_{\ell}$ in $\mathtt{GM}$
9:         Vector core waits for Cube core
11:         for each $s$-tile of $\bm{y}_{\ell}$: $\bm{y}_{s}$ do
12:              $\bm{y}_{s}\leftarrow\bm{y}_{s}+\textit{partial}$triangleright In-place add
[…]
14:         end for
15:         Copy $\bm{y}_{\ell}$ from $\mathtt{UB}$ to $\mathtt{GM}$
16:     end for
17:     Return $\bm{y}$
18:end procedure
Algorithm 1 Scan Cube-Vector (ScanU)
The second algorithm, ScanUL1 (Algorithm 2), is an Ascend adaptation of (Dakkak et al., 2019, Algorithm 6) and is based on a matrix identity that expresses the scan of an array $\bm{z}$ of length $\ell$ using matrix operations. View $\bm{z}$ as a square row-major matrix $\bm{A}$ of size $s=\lceil\sqrt{\ell}\rceil$ (pad with zeroes if needed). Given $\bm{z}$, the inclusive scan of $\bm{z}$ ($\texttt{scan}(\bm{z})$) can be computed as:
(1) $\texttt{scan}(\bm{z})=\bm{A}_{s}\ @\ \bm{U}_{s}+\bm{L}^{-}_{s}\ @\ \bm{A}_{s}\ @\ \bm{1}_{s},$
ignoring any padded values. Equation 1 first appeared in (Dakkak et al., 2019). ScanUL1 uses Equation 1 to scan each consecutive tile of size $\ell$ (Lines 6-12 of Algorithm 2), and then propagates the last value of the partial sums sequentially. In a high-level, for each tile of size $\ell$, the cube unit evaluates Equation 1 with the following sequence of matrix operations:
[…]
$\bm{C}_{2}$ $=$ $\bm{A}_{s}\ @\ \bm{U}_{s}$
$\bm{C}_{2}$ $=$ $\bm{C}_{2}+\bm{L}_{s}^{-}\ @\ \bm{C}_{1}.$
The above sequence of matrix operations has two advantageous properties with respect to data movements. The first two steps of the above sequence share the left matrix operand $\bm{A}$, allowing us to load $\bm{A}$ only once in $\mathtt{L0A}$. Moreover, the third step effectively utilizes the accumulation buffer of the cube unit since $\bm{C}_{2}$ is reused in the last two steps. Once the local scan of a tile of size $\ell$ is computed, a single vector core adds the last value of the previous scanned tile to the current tile (see Lines $14-16$ of Algorithm 2).
1:procedure ScanUL1($\bm{x}$, $s$)
2:     Let $\bm{y}$ be the output array
3:     $\textit{partial}\leftarrow 0$triangleright Accumulation value (Vector Unit)
4:     Load $\bm{U}_{s},\bm{L}_{s}^{-},\bm{1}_{s}$ in $\mathtt{L1}$
5:     for each $s^{2}$-tile of $\bm{x}$: $\bm{x}_{\ell}$ do $\triangleright$ Pipelined execution
6:         Load $\bm{x}_{\ell}$ to $\mathtt{L0A}$ and $\bm{1}_{s}$ to $\mathtt{L0B}$
[…]
8:         Copy $\bm{C}_{1}$ from $\mathtt{L0C}$ to $\mathtt{L1}$
9:         Load $\bm{U}_{s}$ to $\mathtt{L0B}$
10:         $\bm{C}_{2}\leftarrow\bm{A}_{s}\ @\ \bm{U}_{s}$triangleright acc. off, no free inputs
11:         Load $\bm{L}_{s}^{-}$ in $\mathtt{L0A}$ and $\bm{C}_{1}$ in $\mathtt{L0B}$
12:         $\bm{C}_{2}\leftarrow\bm{C}_{2}+\bm{L}_{s}^{-}\ @\ \bm{C}_{1}$triangleright acc. on, free all buffers
13:         Copy $\bm{C}_{2}$ from $\mathtt{L0C}$ to $\bm{y}_{\ell}$ in $\mathtt{GM}$
14:         Vector unit waits for cube unit
15:         Copy $\bm{y}_{\ell}$ from $\mathtt{GM}$ to $\mathtt{UB}$triangleright Vector unit
16:         $\bm{y}_{\ell}\leftarrow$ $\bm{y}_{\ell}+\textit{partial}$
[…]
18:         Copy $\bm{y}_{\ell}$ from $\mathtt{UB}$ to $\mathtt{GM}$
19:     end for
20:     Return $\bm{y}$
21:end procedure
Algorithm 2 ScanUL1 is an Ascend adaptation of (Dakkak et al., 2019, Algorithm 6)
Does cube utilization imply performance?
Although the above scan algorithms demonstrate that it is possible to utilize the cube units for scan, it is unclear if cube utilization translates to performance improvements compared to “vector only” scan algorithms. We provide an experimental evaluation demonstrating the benefits of using the Cube unit. We developed a vector-only kernel that uses the CumSum AscendC API [^1] CumSum API documentation available at https://www.hiascend.com (accessed on $25$ August $2024$). with CumSumInfo parameters set to $128$ and $128$. We also set $s=128$ on the cube scan algorithms to ensure a fair comparison. Figure 3 compares the cube scan algorithms and the vector-only algorithm provided by AscendC. The figure demonstrates a significant performance improvement (5$\times$ for ScanU, and 9.6$\times$ for ScanUL1) compared to the vector-only CumSum algorithm. Moreover, the figure shows that the ScanUL1 scan algorithm has roughly a 2×2$\times$2 × speedup compared to ScanU. The critical insight here is that a more sophisticated usage of the computational capabilities of the cube unit can deliver further significant performance improvements.
Figure 3. Execution time of CumSum AscendC API (vec_only) versus ScanU and ScanUL1 (log-log scale).
[…]
[Execution time of CumSum.]Execution time of CumSum AscendC API (vec_only) versus ScanU and ScanUL1 (log-log scale).
4.2. Multiple (Batched) Scans
Figure 4. Batched scan algorithm based on ScanU. Blue and green denote the input and output array, respectively.
[…]
[Batched scan algorithm based on ScanU.]Batched scan algorithm based on ScanU. Blue and green denote the input and output array, respectively.

### v2

@@ 4.1 Warm-up: single cube scans
Figure 4.2: Execution time of AscendC::CumSum (vec_only) versus ScanU and ScanUL1 (log-log scale).
Work/Span Analysis of Single Core Scans.
We bound the work and span of the single-core algorithms. For work, we separately report the number of square matrix multiplications of size $s$ and the sum of lengths of the vector instructions similar to the Vector-RAM model [8]. The number of matrix multiplications/accumulations for ScanU  and ScanUL1 are $\lceil n/s^{2}\rceil$ and $3\lceil n/s^{2}\rceil$, respectively. The total vector work for both algorithms is $\mathcal{O}(n)$.
Next, we bound the span of the algorithms. The number of vector instructions called by the algorithms ScanU and ScanUL1 are $\mathcal{O}(n/s)$ and $\mathcal{O}(n/s^{2})$, respectively. Notice that the vector length of the ScanUL1 vector operations is $\mathcal{O}(s^{2})$. Since the vector instructions are on the critical path, the span of ScanU and ScanUL1 is $\mathcal{O}(n/s)$ and $\mathcal{O}(n/s^{2})$, respectively. In summary, ScanUL1 has a $\mathcal{O}(s)$-depth speed-up compared to ScanU, but, in practice, the achievable speed-up is $\approx 1.92$ for $s=128$, see Figure 4.2.
4.2 Multi-core Scan
This section presents a multi-core scan algorithm (MCScanAlgorithm 3) to compute the prefix sum of large input arrays. For the sake of clarity, the presented algorithm employs a 1-to-1 vector-to-cube core ratio. Our implementation takes advantage of the 2-to-1 ratio of 910B, but we consider it an implementation detail.
MCScan is similar to the Scan-Scan-Add (SSA) paradigm of computing scans using a hierarchical partition of the input into blocks (top-level of the hierarchy) and tiles within each block, as discussed in Section 2.1. MCScan consists of two phases separated by a global synchronization barrier across all cores (blocks). In contrast to previous methods, MCScan performs partial re-computation of the block-level reductions during its first phase, as we will explain next.
In the first phase, the cube and vector units work in parallel to partially compute the first scan part of SSA simultaneously. The cube units compute the local prefix sums of all consecutive (row) tiles of size $s$ and write them back to global memory. In parallel with the cube units, the vector units compute the reduction over the tiles and then hierarchically reduce the tile reductions on a block-level granularity. The result of these reductions is written in an array $\bm{r}$ where the $i$-th entry equals the reduction of all the values of the $i$-th block. By definition, $\bm{r}$ has length equal to the number of blocks, $B$.
1:procedure MCScan($\bm{x}$, $s$, $B$) $\triangleright$ $B$: number of blocks
2:  Let $\bm{y}$ be the output array
3:  Let $\bm{r}$ be an array of length $B$ in $\mathtt{GM}$
[…]
5:   Load $\bm{U}_{s}$ in $\mathtt{L0B}$ $\triangleright$ Cube Units
6:   for each $\ell$-tile of $\bm{x}[i]$: $\bm{x}_{\ell}$ do
7:     Load $\bm{x}_{\ell}$ from $\mathtt{GM}$ to $\mathtt{L0A}$
8:     $\bm{C}\leftarrow\bm{A}_{s}\ @\ \bm{U}_{s}$
9:     Copy $\bm{C}$ in $\bm{y}[i]$ in $\mathtt{GM}$
10:   Load $\bm{x}[i]$ to $\mathtt{UB}$ $\triangleright$ Vector Units
11:   $r_{i}\leftarrow$ ReduceSum($\bm{x}[i]$)
12:   Write $r_{i}$ on $i$-th entry of $\bm{r}$ in $\mathtt{GM}$
13:  end parfor
14:  SyncAll: Synchronize all cube/vector cores
15:  parfor $i$-th block of $\bm{y}$: $\bm{y}[i]$ do $\triangleright$ Phase II
16:   Load $\bm{r}$ from $\mathtt{GM}$ to $\mathtt{UB}$
17:   $\textit{partial}\leftarrow$ Sum first $i$ entries of $\bm{r}$
18:   for each $\ell$-tile of $\bm{y}[i]$: $\bm{y}_{\ell}$ do
19:     for each $s$-tile of $\bm{y}_{\ell}$: $\bm{y}_{s}$ do
20:      $\bm{y}_{s}\leftarrow\bm{y}_{s}+\textit{partial}$
21:      $\textit{partial}\leftarrow$ last entry of $\bm{y}_{s}$
22:     Copy $\bm{y}_{\ell}$ from $\mathtt{UB}$ to $\mathtt{GM}$
23:  end parfor
24:  Return $\bm{y}$
Algorithm 3 Multi-core Scan (MCScan)
Work/Span Analysis of MCScan.
First, we bound the work of MCScan in terms of the number of matrix multiplications of size $s$, and the number of vector instructions (scalar-vector addition) of length $s$. The number of matrix multiplications of MCScan is $\lceil n/s^{2}\rceil$. The vector instructions of length $s$ consist of $\mathcal{O}(n/s)$ reductions in the first phase, and $\mathcal{O}(n/s)$ scalar-vector additions in the second phase. Hence, the total number of vector operations is $\mathcal{O}(n/s)$. The critical path of MCScan is on the second phase where the vector units propagate the partial sums, thus the span of MCScan is $\mathcal{O}(\frac{n}{sB})$.
Exclusive scan and int8 support.
We have added support for exclusive scans and input with int8 data type to our MCScan. The interested reader is referred to Appendix B.2 for more details.
5 Operators based on Scan
In this section, we revisit several computational parallel primitives based on parallel scans [6, 44]. Scan-based primitive include weighted sampling, split and compress/compact (equivalent to the masked_select Pytorch operator). It is well-known that radix sort can be implemented on top of split [7, Section 1.3]. Also top-k can be implemented on top of split using a partial quick-sort/select approach [9].
Interestingly enough, current AI workloads like Large Language Model (LLM) inference make implicitly heavy use of scan-based computational primitives, including top-$k$ and top-$p$ (nucleus) sampling [18], see also [16]. The top-$p$ sampling implementation of the popular open-source model Llama3 [49] contains a batched sorting and prefix sum operation as the first two PyTorch operations, see [39].
Figure 5.1: A diagram of well-known parallel scan applications considered here along with their dependencies.
In the rest of the section, we describe the split, compress, radix sort, top-$k$, top-$p$ sampling and weighted sampling operators in more detail. Figure 5.1 depicts the dependencies between the scan-based primitives. Top-$p$ sampling extensively uses multiple scan operations, as discussed in Section 5.5.
5.1 Split
5.2 Compress
5.3 Radix Sort
Radix sort is a well-known application of the split operator [7, 10]. A radix sort algorithm loops over the bits of the input elements, starting at the least significant bit and executes a split where the mask is obtained by reading the corresponding bit (radix) on each iteration. We implement a Least-Significant Bit (LSB) radix sort in AscendC using the split operator based on the MCScan algorithm. We implemented an additional vector-only kernel, RadixSingle, that extracts the radices of the inputs before the execution of the split. RadixSingle makes use of the AscendC vector instructions ShiftRight and Not to create the input mask for split. Additional pre-processing and post-processing phases are needed to support floats; see Exercises 8 and 9 in [26, Section 5.2.5]. The pre-processing phase encodes all the input elements by inverting the Most Significant Bit (MSB) of positive numbers and all the bits of the negative numbers. Applying an unsigned integer radix sort on the encoded elements will correctly order them. The post-processing phase is needed to decode the elements back to the original value. We have implemented the pre- and post-processing steps using AscendC bit-wise vector instructions, and thus, support sorting of fp16.
The paper on evaluating radix sort using the Connection Machine (CM-2) came to our attention in the later stages of our radix sort development [10]. In this work, the authors share several important implementation details for radix sort and floats that are replicated here.
5.4 Top-$k$
Top-$k$ selection is an essential operation in various settings, including similarity search queries [23] and Large Language Models (LLMs) inference where the output tokens are typically sampled from the $k$ tokens having the highest probability for the given context [28]. The interested reader is referred to a recent survey on parallel top-$k$ [51]. A recent work on top-$k$ is Radik, a Radix-based GPU implementation that scales well for large values of $k$ [31, 32].
5.5 Top-$p$ or Nucleus Sampling
Top-$p$ sampling in Large Language Model inference is an additional operation that applies sort and scan on the token probability vector [18]. These operations are usually batched with a constant batch size. Interestingly, if the sorting step is implemented using radix sort, the top-$p$ sampling operator becomes a scan-intensive operator! Indeed, top-$p$ executes 17 scans for each batch: 16 scan operations for radix sort (one scan per bit, fp16) and an additional scan required by the algorithm.
Due to space constraints, the study of the weighted sampling operator is deferred to the Appendix B.3.
6 Experimental Evaluation
We evaluate the multi-core scan, compress, radix sort, batch scan, and top-$p$ sampling. We have implemented all proposed algorithms in C++17 using the AscendC programming framework discussed in Section A. We used the Ascend CANN toolkit 8.0.RC3.alpha002 with Ascend firmware and drivers versions 1.0 and 23.0.0, respectively. All evaluations are performed on Huawei’s Ascend 910B4 accelerator. In particular, 910B4 contains 20 Cube Units and 40 Vector Units (the vector-to-cube units ratio is 2-to-1). The theoretical memory bandwidth of 910B4 is $800$ GB/s. The host CPU is an AMD EPYC processor running on Ubuntu 22.04.
All timing measurements are collected using the PyTorch profiler functionality in Python. We used Ascend’s PyTorch adapter [^2] https://gitee.com/ascend/pytorch with version v2.1.0 to report all our PyTorch-related measurements. To wrap our custom AscendC operators in PyTorch, we used the open-sourced operator plugin framework at https://gitee.com/ascend/op-plugin. op-plugin allows its users to easily define a custom PyTorch operator. Before wrapping our operators using the op-plugin, we used the msopgen [^3] https://www.hiascend.com/en/document tool to wrap the AscendC operators.
Evaluation against other accelerators/architectures
6.1 Multi-core Scan (MCScan)
Figure 6.1: Bandwidth of MCScan (Algorithm 3) for $s=32,64,128$. MCScan has $15.2\times$ speedup against ScanU on 910B4 (20 AI cores).
Figure 6.1 depicts the performance of MCScan (Algorithm 3) on Ascend 910B4 versus the state-of-the-art (torch.cumsum). We integrated the multi-core algorithm in PyTorch to ensure a fair comparison and expose it as a custom PyTorch operator. The custom PyTorch operator statically pre-allocates an upper triangular all-ones matrix $\bm{U}_{s}$ for all $s=32,64,128$. The baseline operator doesn’t use the cube unit, while MCScan takes advantage of all the computing units reaching up to 37.5% of theoretical memory bandwidth (peak bandwidth is 800GB/s for 910B4). To have a more solid evaluation of our implementation, we compare it to a copy kernel that performs a memory copy; we used the torch.clone(). A clear trend is that the larger the matrix multiplication dimension $s$ is, the better the performance of the multi-core scan. $s=128$ (almost) maximizes the utilization of the level-0 scratchpad memories $\mathtt{L0A}$ and $\mathtt{L0B}$ of the cube unit. We foresee that increasing the matrix multiplication tile size could lead to further performance improvements, but we leave this investigation to future work.
In addition, we investigate the additional performance benefits of taking advantage of the lower-precision input data (int8) capability of the cube unit; the interested reader is deferred to Section B.4.
6.1.1 Scan Improvements
Figure 6.2: MCScan versus multi-core ScanUL1
We have implemented two optimizations for MCScan. First, we implemented an L2 cache optimization that we call “L2 cache splitting”. This L2 cache optimization splits the input into consecutive chunks so that the input and intermediate arrays fit in the L2 cache during the processing of each chunk. Each chunk is processed serially, and the last value of each chunk is propagated to the next chunk to carry ahead the partial prefix sums. Figure 6.3 shows a performance improvement of 25% for MCScan when the “L2 cache splitting” optimization is enabled. This optimization is already integrated in the MCScan.
Figure 6.3: L2 cache optimization of MCScan ($s=128$).
Second, we designed and implemented a multi-core ScanUL1 algorithm for fp16 by replacing the single matrix multiplication during Phase I of MCScan with the three matrix multiplications of ScanUL1. We call the resulting algorithm, MCScanUL1Figure 6.2 shows a performance improvement of 33% of MCScanUL1 versus MCScan and MCScanUL1 reaches up to 74.9% of the memcopy memory bandwidth as advertised in the abstract. The improvement comes from the fact that ScanUL1 has an $\mathcal{O}(s)$-depth speedup compared to ScanU, see work/span discussion in Section 4.1.
In addition, we provide a performance breakdown analysis of the up-sweep (Phase I) and down-sweep (Phase II) phases of both MCScan and MCScanUL1 in Figure 6.4. The figure shows that, although the up-sweep phase of MCScanUL1 is slightly longer than the up-sweep phase of MCScan, the down-sweep phase of MCScanUL1 is significantly shorter.
Figure 6.4: Break-down of up/down sweeps of scans.
6.2 Compress
Figure 6.5: Bandwidth of compress operator based on MCScan ($s=32,64,128$) against torch.masked_select.
Figure 6.5 depicts the performance comparison between compress versus the baseline PyTorch masked_select. We set the mask so that each mask entry is independently set to true or false uniformly at random. The figure indicates that the baseline masked_select operator is not optimized on Ascend, and a code investigation reveals that the baseline does not use the vector or cube units. On the other hand, our Compress kernel reaches up to 160GB/s (20% of peak memory bandwidth).
6.3 Radix Sort
We modified the radix sort operator to additionally return indices that correspond to the input index of each output element. This modification ensures a fair comparison with the sort operator provided by the PyTorch Ascend adapter. The modification is based on the SplitInd operator by keeping track of the output indices on each split application. Our radix sort implementation is stable and supports unsigned (or signed) integers and floats (fp1616).
Figure 6.6: Comparison between radix sort and baseline torch_npu.sort() for half-precision (fp1616).
Figure 6.6 shows the performance of a parallel fp16 radix sort implementation using MCScan with input data type int8 (Algorithm 3) to perform the parallel split step. For input lengths greater than 525K, our “textbook” implementation of radix sort delivers a speedup between $1.3\times$ and $3.3\times$ compared to the torch.sort() baseline.
Performance Opportunities for small bit-width inputs
We expect additional performance improvements for low-precision data types (low bit-width) since the number of radix sort iterations equals the input bit-width. Indeed, the trend in AI accelerators is to introduce low-precision formats like $8$-bit floats or $4$-bit integers (int4) [42]. Therefore, an additional performance improvement ($2\times$) for radix sorting in low-precision $8$-bit scenarios is expected without any extra development effort, see Figure B.4.
6.4 Batched Scan
Figure 6.7 depicts the performance of the batched scan kernels on Ascend for varying input batch sizes and input lengths equal to $65K$. We depict the memory bandwidth achieved for tiling parameters $s=16,32,64$ and $128$. Our proposed batch scan operators for $s=64$ and $128$ reach up to $400$ GB/s. Interestingly enough, for smaller values of $s=16,32$, the performance of the proposed batch scan kernels is poor. In addition, the performance of our proposed batch scan kernel for $s=16$ and the baseline is similar.
Figure 6.7: Bandwidth of batched scan based on Algorithm 1 for increasing batch sizes and $s=16,32,64,128$. Input length is $65K$.
6.5 Top-$p$ Sampling
Figure 6.8: Execution time (in milliseconds) of top-$p$ sampling of Llama3 operator for a single batch. The baseline is labeled PyTorch and supports only discrete distributions with at most $2^{24}$ elements.
Figure 6.8 depicts the execution time of drawing one sample using the top-$p$ sampler as it is implemented in the Llama3 model, see [39]. PyTorch corresponds to the Ascend implementation using the baseline sort and cumsum operators. The lineplots with labels $s=32,64,128$ are similar to the baseline by replacing the sort() and cumsum() operators with the proposed radix sort and multi-core scan, respectively. The figure shows that the baseline top-$p$ sampling implementation scales poorly, mainly because the baseline torch.cumsum operator is not optimized for Ascend.
7 Conclusion
We developed and evaluated efficient parallel scan algorithms tailored to the Ascend architecture by leveraging the power of the cube (matrix multiplication) unit. Our results demonstrate substantial performance improvements, with speedups ranging from $5\times$ to $9.6\times$ compared to vector-only implementations for sufficiently large input lengths. Additionally, we presented a multi-core Ascend scan algorithm that fully utilizes both the cube and vector units of Ascend, reaching up to 74.9% of the theoretical memory bandwidth.
Last, we extended our contributions to include crucial computational kernels for AI workloads, such as parallel split, compress/compact, and top-$p$ (nucleus) sampling, all exhibiting significant performance gains. Furthermore, our optimized implementation of radix sort, which utilizes matrix multiplications for parallel splits, showcases the potential of matrix engines in enhancing complex operations, offering up to $3.3\times$ speedup over the baseline.
A AscendC Programming Model
@@
B Additional Material for Scans
B.1 Multiple (Batched) Scans